### v2

@@ 5.1.1 Approximation on the Unit Cube
The starting point of our construction is the explicit representation of B-spline functions. According to [41, Theorem 4.14 and (4.49)], the B-splines $B_{i}^{k}(x)$ can be expressed as:
@@ 5.2 Lower Bounds
The first term, $\mathrm{ReLU}^{k-2}(\cdot)$, belongs to the class of activations for $\mathcal{F}^{k-2,k-1}$. By the induction hypothesis, the second term $\sum_{j}w_{j}D_{i}u_{j}^{L}(x)$, can be realized by a network in $\mathcal{F}^{k-2,k-1}$. By Lemma 2, the product of these two functions can be constructed by a new $\mathrm{ReLU}^{k-1}$network. Notice that these two factors belong to the $\mathcal{F}^{k-2,k-1}$ network class with respective widths of $W$ and $(Q+2)W$, which implies the resulting product network will have a width of $(Q+3)W$. Thus, the claim holds for $s=1$.
@@ A.2 Riemannian Manifolds
    The injectivity radius is uniformly bounded below by a positive constant. That is, there exists a constant $\delta>0$ such that
    $\mathrm{inj}(\mathcal{M}^{d})\geq\delta.$
    This ensures that for any point $x\in\mathcal{M}^{d}$, the exponential map $\exp_{x}$ is a diffeomorphism from the ball $B(0,\delta)\subset T_{x}\mathcal{M}^{d}$ to its image in $\mathcal{M}^{d}$.
@@
    The components of the metric tensor and all their derivatives are uniformly bounded in normal coordinates. Specifically, there exist constants $C>0$ and $\{C_{\alpha}\}_{\alpha}$ (independent of the choice of coordinates) such that in any normal coordinate chart, the following hold for the metric components $g_{ij}$:
    $\det(g_{ij})\geq C\quad\text{and}\quad|\partial^{\alpha}g_{ij}|\leq C_{\alpha}$
    for all multi-indices $\alpha$.
@@ B.1 Tensor Product Splines
The proof can be found in [41, Theorem 12.5 and Theorem 12.6].
@@
This proof follows directly from [41, Theorem 4.22].

### v3

@@ 6.1.1 Approximation on the Unit Cube
The starting point of our construction is the explicit representation of B-spline functions. According to [44, Theorem 4.14 and (4.49)], the B-splines $B_{i}^{k}(x)$ can be expressed as:
@@ 6.2 Lower Bounds
The first term, $\mathrm{ReLU}^{k-2}(\cdot)$, belongs to the class of activations for $\mathcal{F}^{k-2,k-1}$. By the induction hypothesis, the second term $\sum_{j}w_{j}D_{i}u_{j}^{L}(x)$, can be realized by a network in $\mathcal{F}^{k-2,k-1}$. By Lemma 2, the product of these two functions can be constructed by a new $\mathrm{ReLU}^{k-1}$network. Notice that these two factors belong to the $\mathcal{F}^{k-2,k-1}$ network class with respective widths of $W$ and $(L+2)W$, which implies the resulting product network will have a width of $(L+3)W$. Thus, the claim holds for $s=1$.
@@ A.2 Riemannian Manifolds
    The injectivity radius is uniformly bounded below by a positive constant. That is, there exists a constant $\delta>0$ such that inj(M^d) δ. This ensures that for any point $x\in\mathcal{M}^{d}$, the exponential map $\exp_{x}$ is a diffeomorphism from the ball $B(0,\delta)\subset T_{x}\mathcal{M}^{d}$ to its image in $\mathcal{M}^{d}$.
@@
    The components of the metric tensor and all their derivatives are uniformly bounded in normal coordinates. Specifically, there exist constants $C>0$ and $\{C_{\alpha}\}_{\alpha}$ (independent of the choice of coordinates) such that in any normal coordinate chart, the following hold for the metric components $g_{ij}$: det(g_ij) C  and  —^αg_ij— C_α for all multi-indices $\alpha$.
@@ B.1 Tensor Product Splines
The proof can be found in [44, Theorem 12.5 and Theorem 12.6].
@@
This proof follows directly from [44, Theorem 4.22].
@@ B.3 Averaged Taylor Polynomials
Appendix C Laplace–Beltrami Operator on Spheres and the Torus
This section collects basic formulae and spectral descriptions of the Laplace–Beltrami operator on the unit sphere and on an embedded torus. See [14] for further details on spherical harmonics.
C.1 Sphere Laplacian
Let the unit sphere be $\mathbb{S}^{d-1}=\{x\in\mathbb{R}^{d}:\|x\|=1\}$. For any smooth $f:\mathbb{S}^{d-1}\to\mathbb{R}$ take a smooth extension $\tilde{f}$ to a neighborhood in $\mathbb{R}^{d}$ (for instance the 0-homogeneous extension $\tilde{f}(x)=f(x/\|x\|)$). At $x\in\mathbb{S}^{d-1}$ define the tangential projection matrix
$P_{x}:=I-xx^{\top}\in\mathbb{R}^{d\times d},$
so that the surface (tangential) gradient can be written as
$\nabla_{\mathbb{S}^{d-1}}f(x)=P_{x}\,\nabla_{\mathbb{R}^{d}}\tilde{f}(x),\qquad|x|=1.$
The Laplace–Beltrami operator on the sphere is the tangential divergence of the tangential gradient:
$\Delta_{\mathbb{S}^{d-1}}f(x)=\operatorname{div}_{\mathbb{R}^{d}}\!\big(P_{x}\,\nabla_{\mathbb{R}^{d}}\tilde{f}\big)(x),\qquad x\in\mathbb{S}^{d-1},$
and this representation is independent of the chosen extension $\tilde{f}$. Expanding in ambient coordinates yields the angular Laplacian
$\Delta_{0}=\sum_{i=1}^{d}\frac{\partial^{2}}{\partial x_{i}^{2}}-\sum_{i=1}^{d}\sum_{j=1}^{d}x_{i}x_{j}\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}-(d-1)\sum_{i=1}^{d}x_{i}\frac{\partial}{\partial x_{i}}.$
Introducing angular derivatives
$D_{i,j}=x_{i}\frac{\partial}{\partial x_{j}}-x_{j}\frac{\partial}{\partial x_{i}},\qquad 1\leq i<j\leq d,$
one has the angular representation
$\Delta_{\mathbb{S}^{d-1}}=\sum_{1\leq i<j\leq d}D_{i,j}^{2}.$
Spherical harmonics are defined spectrally as eigenfunctions of $\Delta_{\mathbb{S}^{d-1}}$. For each integer $n\geq 0$ the eigenvalue is
$\lambda_{n}=n(n+d-2),$
and the corresponding eigenspace is given by the restrictions to the sphere of homogeneous harmonic polynomials of degree $n$; its multiplicity is
$m(n)=\binom{n+d-1}{n}-\binom{n+d-3}{n-2}$.
In $L^{2}(\mathbb{S}^{d-1})$ choose an orthonormal basis $\{Y_{n,k}\}_{n\geq 0,\;1\leq k\leq m(n)}$ so that
$\Delta_{\mathbb{S}^{d-1}}Y_{n,k}=-\lambda_{n}\,Y_{n,k}.$
Any $f\in L^{2}(\mathbb{S}^{d-1})$ admits the expansion
$f(\theta)=\sum_{n=0}^{\infty}\sum_{k=1}^{m(n)}f_{n,k}\,Y_{n,k}(\theta),\qquad f_{n,k}=\int_{\mathbb{S}^{d-1}}f(\theta)\,\overline{Y_{n,k}(\theta)}\,\mathrm{d}\omega(\theta),$
with Parseval’s identity
$\|f\|_{L^{2}(\mathbb{S}^{d-1})}^{2}=\sum_{n=0}^{\infty}\sum_{k=1}^{m(n)}|f_{n,k}|^{2}.$
From the spectral representation one obtains the energy identities
$\|\nabla_{\mathbb{S}^{d-1}}f\|_{L^{2}}^{2}=\sum_{n,k}\lambda_{n}\,|f_{n,k}|^{2},\qquad\|\Delta_{\mathbb{S}^{d-1}}f\|_{L^{2}}^{2}=\sum_{n,k}\lambda_{n}^{2}\,|f_{n,k}|^{2},$
and hence the Sobolev $H^{s}(\mathbb{S}^{d-1})$ norm admits the spectral weight representation
$\|f\|_{H^{s}(\mathbb{S}^{d-1})}^{2}\simeq\sum_{n=0}^{\infty}(1+\lambda_{n})^{s}\sum_{k=1}^{m(n)}|f_{n,k}|^{2}.$
C.2 Torus Laplacian
We consider the embedded torus
$\mathbb{T}^{2}_{R,r}=X_{R,r}([0,2\pi)^{2})\subset\mathbb{R}^{3},\qquad X_{R,r}(u,v)=\big((R+r\cos v)\cos u,\;(R+r\cos v)\sin u,\;r\sin v\big),$
with $R>r>0$. The first fundamental form induced by this parametrization is diagonal,
$g=\begin{pmatrix}(R+r\cos v)^{2}&0\\[4.0pt] 0&r^{2}\end{pmatrix}$,
so the area element is $\sqrt{|g|}=r\,(R+r\cos v)$. The inverse metric components are $g^{uu}=(R+r\cos v)^{-2}$, $g^{vv}=r^{-2}$. Therefore the Laplace–Beltrami operator in the coordinates $(u,v)$ takes the divergence form
$\Delta_{\mathbb{T}^{2}_{R,r}}f=\frac{1}{\sqrt{|g|}}\Big[\partial_{u}\big(\sqrt{|g|}\,g^{uu}\,\partial_{u}f\big)+\partial_{v}\big(\sqrt{|g|}\,g^{vv}\,\partial_{v}f\big)\Big],$ (C.1)
or equivalently
$\Delta_{\mathbb{T}^{2}_{R,r}}f=\frac{1}{r(R+r\cos v)}\Big\{\partial_{u}\!\Big(\frac{r}{R+r\cos v}\partial_{u}f\Big)+\partial_{v}\!\Big(\frac{R+r\cos v}{r}\partial_{v}f\Big)\Big\}.$
For the analytic Fourier-type family used in the experiments one obtains closed-form expressions. If
$g(u,v)=A\cos(mu)\qquad(m\in\mathbb{Z}_{\geq 0}),$
then $g$ is independent of $v$ and from (C.1) one gets
$\Delta_{\mathbb{T}^{2}_{R,r}}[A\cos(mu)]=-\frac{m^{2}\,A\cos(mu)}{(R+r\cos v)^{2}},$ (C.2)
which is explicit (the right-hand side depends on $v$ through the metric factor). For a $v$-mode
$h(u,v)=B\sin(nv)\qquad(n\in\mathbb{Z}_{\geq 0}),$
a direct computation yields the closed form
$\Delta_{\mathbb{T}^{2}_{R,r}}[B\sin(nv)]=\frac{B\,n}{r^{2}(R+r\cos v)}\Big[-r\sin v\cos(nv)-n(R+r\cos v)\sin(nv)\Big],$ (C.3)
and analogous expressions hold for $\cos(nv)$ and linear combinations thereof. These closed-form images make $\Delta_{\mathbb{T}^{2}_{R,r}}g$ available exactly for supervision in the loss.
The formulae above provide the pointwise and spectral tools used in the numerical experiments: for the sphere one uses the ambient-space/angular representations and spherical-harmonic expansion, while for the torus one evaluates $\Delta_{\mathbb{T}^{2}_{R,r}}$ via (C.1) and uses (C.2)–(C.3) for the analytic Fourier targets.